A nursery sells Kentucky Blue Grass seed for $5.75 per pound and Tall Fescue for $4.50 a pound. The nursery sells a mixture of the two kinds of seed for $5.25 per pound. How much Tall Fescue does the nursery need to use in 5 pounds of the mixture.
Answer by Boreal(15235)

KBG=x
TF=y
x+y=5
5.75x+4.50y=5.25*5=26.125
multiply the top by -4.5 to eliminate the y
-4.5x-4.5y=-22.5
Now add the two
1.25x=3.625
x=2.9 pounds KBG@5.75=$16.675
2.1 pounds TF@4.5=$9.45
They add to $26.125.